3. A rubber chicken is launched straight up at speed v from ground level. Find each of the following if the launch speed is tripled (in terms of any constants and v).
max height
hang time
impact speed
1
Expert's answer
2021-10-03T14:13:53-0400

Solution.

"h_{max}=\dfrac{v^2}{2g};"


"h_{max}=\dfrac{9v^2}{2g};"

"t= 2\sdot \dfrac{v}{g};"

"t=2\sdot \dfrac{3v}{g};"

"v_{impact}=v;"

"v_{impact}=3v;"

Answer: "h_{max}=\dfrac{9v^2}{2g};"

"t=2\sdot \dfrac{3v}{g};"

"v_{impact}=3v;"